# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v7

@@ Abstract
Current artificial intelligence systems continue to fall short of human-like general intelligence. Existing evaluation frameworks, which focus on language or perception tasks, fail to capture the essence of generality or provide actionable guidance. The Artificial General Intelligence Testbed (AGITB) introduces a novel benchmarking suite comprising fourteen elementary tests, thirteen of which are fully automated.
AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalisation—that parallel principles of biological information processing. Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning. The human cortex satisfies all tests, whereas no current AI system meets the full AGITB criteria, demonstrating its value as a rigorous, interpretable, and actionable benchmark for evaluating progress toward artificial general intelligence. A reference implementation of AGITB is freely available on GitHub.
@@ 1 Introduction
Despite rapid advances in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence. Marcus2020 observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. Surface-level competence obscures a deeper lack of understanding, which is an essential prerequisite for artificial general intelligence (AGI).
@@
AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, thirteen of which are implemented as fully automatable and domain-agnostic tests. Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour. To position AGITB within the broader landscape of AGI evaluation, we include a comparative analysis with the Abstraction and Reasoning Corpus (ARC; Chollet2019) and the NeuroBench framework (Yik2025).
@@ 2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
@@ 2.1 A Benchmark That Only Humans and AGI Can Meet
In alignment with this biologically grounded perspective, the AGITB departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
@@
Rooted in the view—advanced by Hawkins2004—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory spike trains, rather than disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The central premise is that the capacity to pass the AGITB constitutes a necessary condition for the existence of AGI. Non-AGI systems cannot meet AGITB’s requirements. Such a premise cannot be proven; it can only be rejected by any non-AGI system that meets the AGITB criteria. The fact that humans can pass the AGITB suggests that AGITB requires a solution beyond the current state-of-the-art AI.
Unfortunately, without a clear definition of intelligence, we can only speculate that AGITB defines the necessary (but not sufficient) building blocks for the emergence of genuine, domain-independent intelligence. This assumption of necessity stems from the fact that, except for human brains, no existing artificial system has met its requirements.
However, an AGITB-passing system is not automatically also a fully fledged AGI; after all, it is not required to express higher-level skills, like reasoning or natural language fluency. However, a failure to meet the AGITB’s requirements likely indicates a deficiency in foundational general cognitive capabilities.
3.1 Components
AGITB evaluates the user-supplied AGI model as a black box that predicts future inputs from the historical sequence of observed signals. Each input comprises ten bits, the specific semantics of which are immaterial; each bit may correspond to an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
@@ 3.1 Components
Figure 1: (a) Example of a 10-bit input $x$ with 4 bits set.
(b) Example of an input sequence showing a temporal pattern of period four.
@@ 3.2 Operation
The testbed presents the model with a stream of inputs over time. At each time step $t$, the model receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
Figure 2: Iterative adaptation in discrete time. At the previous time step (t1t-1), the model issued a prediction $x_{t}^{*}$, with the second bit incorrect in this instance. After observing the realized input $x_{t}$, the model updates and subsequently produces a correct one-step-ahead prediction $x_{t+1}^{*}$ for the sequence shown in Figure 1b.
@@ 3.3 The 14 AGI Requirements
Requirement 1 – Unbiased start
Assertion: Models that have received no input are considered empty and thus equal.
Assertion: An empty model predicts an empty input.
AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs. Instead, they learn through interaction with their environment. Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics. While some reflexes may be genetically predetermined, these do not constitute genuine understanding. In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
A model that predicts non-empty (non-zero) inputs prior to receiving data is considered biased, as it encodes unwarranted assumptions about the future. To ensure neutrality, models must initialise with empty predictions and remain unbiased until conditioned on actual input.
Two model instances that have received no input should compare as equal, as neither has been influenced by prior experience. Here, "empty" refers to a lack of informational content, not structural capability. The model must possess an intrinsic organisational architecture capable of learning, even in the absence of prior data.
The test operationalises, in empirical form, the philosophical challenge posed by classical symbolic systems: the emergence of meaning without prior grounding.
Requirement 2 – Bias
Assertion: Any model that has received input can no longer be regarded as unbiased.
@@
Requirement 3 – Determinism
Assertion: If two cortices are equal, they must have received the same inputs.
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing. Similarly, in AGITB, two functionally identical models must have experienced identical input histories. Any difference in input must lead to a divergence in state, reinforcing the principle that behaviour and internal state are fully determined by input alone. The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave2016 in his discussion of free will.
Determinism at the level of neural signal processing remains a necessary condition for the stable and reproducible functioning of the brain, even if its manifestations at the cognitive level are obscured by complexity and individual variation.
Requirement 4 – Sensitivity
Assertion: Two different cortices remain different, even if they experience long exposure to identical inputs.
Cortical systems exhibit chaotic sensitivity to initial conditions. Small differences in early experiences or internal states can lead to divergent trajectories over time. This deterministic sensitivity, amplified through complex interactions, accounts for the illusion of unpredictability in decision-making.
Requirement 5 – Time
Assertion: Changing the input order results in a different model state.
Because the model updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation. The ability to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
However, sensitivity to input order alone is not sufficient. Since temporal coding makes time an intrinsic property of neural data, Requirement 6 is designed to rule out models that perform equally well on temporally invariant data, as such models fail to exploit temporal information.
Requirement 6 – Refractory period
Assertion: The model must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
@@
Biological intelligence depends on discrete spikes for signal transmission and learning. AGITB enforces refractory periods to reflect this constraint, capturing the biological principle that a neuron cannot fire again immediately after activation.
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner2002 showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
The refractory period requirement does not preclude the possibility of more general solutions. A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot simulate the absence of such dynamics.
Requirement 7 – Temporal adaptability
@@
This test assesses the model’s ability to learn across different temporal scales. Unlike rigid temporal pattern-matching systems, intelligent models should be able to detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
Requirement 8 – Stagnation
@@
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test assesses whether such limits emerge over time when the model is exposed to input sequences that would otherwise be learnable.
Requirement 9 – Content sensitivity
@@
The structure and regularity of an input sequence influence the rate at which a model adapts, where adaptation time is defined as the number of iterations required to predict the entire temporal pattern accurately. Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
Requirement 10 – Context sensitivity
@@
Past learning influences how new information is integrated. A model with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganisation.
Requirement 11 – Unobservability
Assertion: Distinct cortices may exhibit the same observable behaviour.
Identical external outputs can arise from different internal states. This many-to-one mapping underscores that observable behaviour alone cannot reveal the underlying structure or history of a model’s internal dynamics.
Requirement 12 – Denoising
@@
An intelligent model should be capable of efficiently memorising input patterns in a manner that enables the recall of previously observed sequences even when these are perturbed by noise. When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks. Average performance is used as the evaluation criterion, since random models may occasionally produce correct predictions by chance without reflecting genuine learning or structural understanding.
Test 12 is designed as a deliberately conservative benchmark. The model is required to outperform random guessing in all 100 independent trials. This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it serves as a stringent robustness check. The design ensures that only large and systematic performance gains result in a pass. Modest or marginal improvements, while potentially real, are intentionally treated as failures, as the objective is to identify only clear and substantial advances in model capability.
@@
Only models capable of generalisation can derive lasting benefit from prior learning. After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs. Average performance is used as the evaluation criterion, since even a random model may occasionally produce correct predictions by chance without demonstrating genuine understanding or the acquisition of structured knowledge.
The generalisation assessment follows the same conservative design as Test 12. By requiring success across all repetitions, the procedure establishes a stringent threshold whereby a single underperforming trial constitutes failure. Consequently, the test exhibits minimal tolerance for noise or marginal effects. This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models demonstrating a clear and consistent advantage achieve a passing result.
Requirement 14 – Bounded Prediction Latency
A model must provide low-latency, signal-level predictions to ensure functional viability. Biological brains achieve approximately constant reaction times through extensive parallel processing, such that per-input latency remains bounded and stable under expected workloads.
No universal automated test executed on serial hardware can feasibly verify this property across all possible model implementations. Consequently, model developers are responsible for ensuring that their architectures permit efficient parallel execution with bounded latency. In this regard, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
@@ 3.4 Search space
To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory. In AGITB, each task typically involves predicting a random temporal sequence of seven binary inputs, each comprising ten bits. This defines a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
The refractory period is a biologically inspired constraint that prohibits any neuron (bit) from firing in consecutive time steps. This restriction substantially reduces the number of valid input sequences. There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits, assuming that a $1$ never carries over to the next time step, where $F_{i}$ is the $i$-th Fibonacci number with $F_{0}=0$.
In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence. The number of distinct cyclic temporal input sequences with the refractory constraint is $|S"|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
The use of length-seven sequences with ten-bit inputs is considered sufficient to detect non-AGI behaviour while maintaining computational efficiency. Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing execution time. The current configuration ensures that each test remains computationally feasible and diagnostically informative.
Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure. However, the sheer size of the search space renders explicit optimisation —or "teaching to the test"—computationally infeasible. Given that real-world sensory inputs may encompass tens of thousands of bits, a genuine AGI system must therefore employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from input data.

### v8

@@ Abstract
Current AI systems demonstrate remarkable capabilities yet remain specialised, in part because no unified measure of general intelligence has been established. Existing evaluation frameworks, which focus primarily on language or perception tasks, offer limited insight into generality. The Artificial General Intelligence Testbed (AGITB) introduces a complementary benchmarking suite of fourteen elementary tests, with thirteen implemented as fully automated procedures.
AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants, such as determinism, sensitivity, and generalisation, that parallel principles of biological information processing. Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning. The human cortex satisfies all tests, whereas no current AI system meets the full AGITB criteria, demonstrating its value as a rigorous, interpretable, and actionable benchmark for evaluating progress toward artificial general intelligence. A reference implementation of AGITB is freely available on GitHub.
@@ 1 Introduction
Despite great achievements in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence. Marcus2020 correctly observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. Surface-level competence obscures a deeper lack of understanding, which is an essential prerequisite for artificial general intelligence (AGI).
@@
AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, with thirteen implemented as fully automated and domain-agnostic tests. Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour. To position AGITB within the broader landscape of AGI evaluation, we include a comparative analysis with the Abstraction and Reasoning Corpus (ARC; Chollet2019) and the NeuroBench framework (Yik2025).
@@ 2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits, such as consciousness, self-awareness, and subjective experience, which remain elusive in artificial systems.
@@ 2.1 A Benchmark That Only Humans and AGI Can Meet
In alignment with this biologically grounded perspective, the AGITB departs from symbolic, high-level evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
@@
Building on the view of Hawkins2004 that intelligence is fundamentally about extracting structure from data to enable prediction, AGITB operates at the level of signal-based forecasting. This approach aligns closely with the functioning of biological intelligence at the cortical level, which processes time-sensitive sensory spike trains rather than disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The central premise of AGITB is that it represents a necessary (though not sufficient) condition for the existence of artificial general intelligence. In the absence of a clear definition of intelligence, this premise cannot be formally proven, though it could be falsified by the existence of a non-AGI system that satisfies all AGITB criteria. The assumption of necessity follows from the fact that, apart from the human brain, no known artificial system has yet fulfilled its requirements, suggesting that a solution lies beyond the current state of the art in artificial intelligence.
AGITB is not, in itself, a sufficient criterion for AGI; a system that passes all tests is not automatically a fully-fledged AGI, as it is not required to exhibit higher-level capabilities such as reasoning or natural language fluency.
3.1 Architecture
AGITB evaluates an AGI model as a black box that predicts the next input based on the historical sequence of observed signals. Each input consists of ten bits, the specific semantics of which are immaterial; each bit may represent an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
@@ 3.1 Architecture
Figure 1: (a) Example of a 10-bit input $x$ with four bits set.
(b) Example of an input sequence.
@@ 3.2 Operation
The testbed presents the AGI model with a stream of inputs over time. At each time step $t$, the model receives an input $x_{t}$ and is required to predict the subsequent input $x_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
Figure 2: Iterative adaptation in discrete time. At the previous time step (t1t-1), the model issued the prediction $x_{t}^{*}$. After observing the realised input $x_{t}$, it adapted its internal state in response to the error in the second bit and subsequently produced the one-step-ahead prediction $x_{t+1}^{*}$.
@@ 3.3 The 14 AGI Requirements
Requirement 1 — Unbiased start
Assertion: A model starts unbiased.
Assertion: An unbiased model produces an empty prediction.
AGITB rests on the foundational assumption that general-purpose learning systems, such as the brain, do not begin with an innate understanding of external inputs but instead acquire meaning through interaction with their environment. Each system must construct semantic content from raw sensory data rather than rely on pre-encoded [^1] Although certain reflexes may be genetically specified, they do not constitute genuine understanding. As Hawkins2004 argue, low-level reflexes are not prerequisites for intelligence but rather evolutionary features of the subcortical “old brain”. knowledge. Accordingly, a model should originate in an unbiased initial state shaped exclusively by subsequent inputs, where “unbiased” denotes the absence of informational content rather than the lack of structural capacity for learning.
A model that produces non-empty predictions prior to receiving any input is considered biased, as it encodes unwarranted assumptions about the future. To ensure neutrality, models must initialise in a completely blank state and remain so until conditioned on actual data. Although a spike-free prediction could be interpreted as a bias toward silence, this convention is justified by biological analogy: neurons do not activate in the absence of input. Two blank model instances should be regarded as equivalent, as neither has been influenced by prior experience.
The unbiased start requirement serves as the foundation for other tests that operationalise the emergence of meaning in the absence of prior grounding.
Requirement 2 — Bias
Assertion: A model that has received input can no longer be regarded as unbiased.
@@
Requirement 3 — Determinism
Assertion: If two models are equivalent, they must have received the same inputs.
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although minor stochastic effects may occur, they do not undermine the rule-governed nature of neural processing. By analogy, AGITB assumes that two functionally identical models must have experienced identical input histories: any difference in input necessarily produces a divergence in state. This principle reinforces the view that a model’s internal state is fully determined by its input history.
Determinism at the level of neural signal processing is necessary for stable, reproducible brain function, whereas the apparent unpredictability of cognition stems from the system’s complexity rather than from genuine indeterminacy (Cave2016).
Requirement 4 — Sensitivity
Assertion: Two distinct models remain distinct, even after prolonged exposure to identical inputs.
Cortical systems exhibit chaotic sensitivity to initial conditions: small variations in early experiences can lead to divergent trajectories over time. This deterministic sensitivity, amplified by complex internal interactions, underlies the apparent unpredictability of decision-making.
Requirement 5 — Time
Assertion: Changing the order of inputs results in a different model state.
Since the model’s state depends on its cumulative history, the sequence of inputs critically shapes learning and adaptation. The capacity to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
However, sensitivity to input order alone is not sufficient. Because temporal coding makes time an intrinsic feature of neural data, Requirement 6 is designed to exclude models that perform equally well on temporally invariant data, as such models fail to make functional use of temporal information.
Requirement 6 — Refractory period
Assertion: The model must be able to adapt to any minimal-period input sequence that respects biologically plausible refractory periods.
@@
Biological intelligence relies on discrete spikes for communication and learning. AGITB enforces refractory periods to reflect this constraint, embodying the biological principle that a neuron cannot fire again immediately after activation.
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner2002 showed that input sequences that lack sufficient variability, such as constant or overly repetitive spiking, fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
This requirement does not exclude the possibility of more general solutions. A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot its absence.
Requirement 7 — Temporal adaptability
@@
This test requires the model to learn and track temporal structure across multiple timescales. Unlike rigid pattern-matching systems, an intelligent model should detect and predict a recurring structure regardless of its exact periodicity. Systems that can accommodate only a single, predetermined temporal scale therefore fail this requirement.
Requirement 8 — Stagnation
@@
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test evaluates whether such a boundary arises when the model is exposed over time to input sequences that would otherwise be learnable.
Requirement 9 — Content sensitivity
@@
The structural complexity of an input sequence influences the rate at which a model adapts, where adaptation time is defined as the number of iterations required for the model to predict the entire temporal pattern accurately. Simple or highly regular sequences typically lead to rapid convergence, whereas irregular or noisy inputs demand longer exposure before the model can reliably capture and reproduce the underlying pattern.
Requirement 10 — Context sensitivity
@@
The model’s current state reflects the cumulative influence of past inputs and therefore provides the context in which new information is processed. When subsequent inputs are consistent with structure already established through prior learning, adaptation proceeds quickly. Conversely, when new inputs conflict with this learned context, the model requires additional time to reorganise its state before accurate prediction becomes possible.
Requirement 11 — Unobservability
Assertion: Distinct models may exhibit the same observable behaviour.
Identical external outputs may arise from distinct internal states. This many-to-one mapping highlights that observable behaviour alone cannot disclose the underlying structure or history of a model’s internal dynamics.
Requirement 12 — Denoising
@@
An intelligent model should be able to recall a previously observed sequence, even when the inputs are perturbed by noise. When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks. Average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
Test 12 is intentionally designed as a conservative benchmark. The model must outperform random guessing in all 100 independent trials. This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it functions as a stringent robustness check. The design ensures that only large, systematic performance gains yield a passing result. Modest or marginal improvements, while potentially real, are intentionally regarded as failures, as the objective is to identify only clear and substantial advances in model capability.
@@
Only models capable of generalisation can derive lasting benefit from prior learning. After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs. As in the preceding requirement, average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
The generalisation assessment follows the same conservative design as Test 12. By requiring success across all repetitions, the procedure imposes a stringent threshold in which a single underperforming trial results in failure. As a consequence, the test tolerates virtually no noise or marginal effects. This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models exhibiting a clear and consistent advantage obtain a passing result.
Requirement 14 — Bounded Prediction Latency
A model must provide low-latency, signal-level predictions to ensure functional viability. Biological brains achieve approximately constant reaction times through extensive parallel processing, keeping per-input latency bounded and stable under expected workloads.
No universal automated test running on serial hardware can feasibly verify this property across all possible model implementations. Accordingly, it is the responsibility of model developers to ensure that their architectures support efficient parallel execution with bounded latency. In this respect, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
@@ 3.4 Search space
To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory. In AGITB, tasks typically involve predicting a random temporal sequence of seven (or more) binary inputs, each comprising ten bits. This yields a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
AGITB incorporates a biologically inspired refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This restriction substantially reduces the number of valid sequences. There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits under the condition that a $1$ never carries over to the next time step, where $F_{i}$ denotes the $i$-th Fibonacci number with $F_{0}=0$.
In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence. The number of distinct cyclic temporal sequences respecting the refractory constraint is $|S^{\prime\prime}|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
The choice of seven-step sequences with ten-bit inputs is sufficient to detect non-AGI behaviour while maintaining computational efficiency. Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing runtime. The current configuration therefore ensures that each test remains both computationally feasible and diagnostically informative.
Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure. However, the sheer size of the search space makes any form of explicit teaching-to-the-test computationally infeasible. Given that real-world sensory inputs may ultimately encompass tens of thousands of bits, a genuine AGI system must employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from high-dimensional data.